Two angles are complementary. The measure of one angle is 38° more than three times the measure of the other. how do i find this? i don't want the answer i just wanna know how to work it out.
Answer by stanbon(75887)

Two angles are complementary. The measure of one angle is 38° more than three times the measure of the other. how do i find this? i don't want the answer i just wanna know how to work it out.
-----------------------
Equations:
x + y = 90
x = 3y + 38
-----------------------
Substitute for "x" and solve for "y"::
3y+38 + y = 90
----
4y = 52
---
y = 13 degrees
x = 90-13 = 77 degrees
-----------------
Cheers,
Stan H.
--------------